# Accurate and thermodynamically consistent hydrogen equation of state for planetary modeling with flow matching

### v1

@@ Acknowledgements.
We acknowledge useful discussions with Lei Wang, Ravit Helled, Gilles Chabrier, and Armin Bergermann. We also acknowledge Cesare Cozza for providing initial data and support with the DFT calculations. H.X. and G.M. acknowledge financial support from the Swiss National Science Foundation (grant PCEFP2_203455). S.H. acknowledges financial support from the Swiss National Science Foundation (grant 200020_215634).References
@@
- Helled et al. [2020] R. Helled, G. Mazzola, and R. Redmer, Understanding dense hydrogen at planetary conditions, Nature Reviews Physics 2, 562 (2020).
@@
- Johnson et al. [1992] T. V. Johnson, C. M. Yeates, and R. Young, Space science reviews volume on Galileo Mission overview, Space Science Reviews 60, 3 (1992).
@@
- Bolton et al. [2017] S. J. Bolton, J. Lunine, D. Stevenson, J. E. P. Connerney, S. Levin, T. C. Owen, F. Bagenal, D. Gautier, A. P. Ingersoll, G. S. Orton, T. Guillot, W. Hubbard, J. Bloxham, A. Coradini, S. K. Stephens, P. Mokashi, R. Thorne, and R. Thorpe, The Juno Mission, Space Science Reviews 213, 5 (2017).
@@
- Miguel, Y. et al. [2016] Miguel, Y., Guillot, T., and Fayon, L., Jupiter internal structure: the effect of different equations of state, A&A 596, A114 (2016).
@@
- Debras and Chabrier [2019] F. Debras and G. Chabrier, New Models of Jupiter in the Context of Juno and Galileo, The Astrophysical Journal 872, 100 (2019).
@@
- Mazevet, S. et al. [2022] Mazevet, S., Licari, A., and Soubiran, F., Benchmarking the ab initio hydrogen equation of state for the interior structure of jupiter, A&A 664, A112 (2022).
@@
- Howard, S. et al. [2023] Howard, S., Guillot, T., Bazot, M., Miguel, Y., Stevenson, D. J., Galanti, E., Kaspi, Y., Hubbard, W. B., Militzer, B., Helled, R., Nettelmann, N., Idini, B., and Bolton, S., Jupiter’s interior from Juno: Equation-of-state uncertainties and dilute core extent, A&A 672, A33 (2023).
@@
- Nettelmann et al. [2021] N. Nettelmann, N. Movshovitz, D. Ni, J. J. Fortney, E. Galanti, Y. Kaspi, R. Helled, C. R. Mankovich, and S. Bolton, Theory of Figures to the Seventh Order and the Interiors of Jupiter and Saturn, The Planetary Science Journal 2, 241 (2021).
@@
- Miguel, Y. et al. [2022] Miguel, Y., Bazot, M., Guillot, T., Howard, S., Galanti, E., Kaspi, Y., Hubbard, W. B., Militzer, B., Helled, R., Atreya, S. K., Connerney, J. E. P., Durante, D., Kulowski, L., Lunine, J. I., Stevenson, D., and Bolton, S., Jupiter’s inhomogeneous envelope, A&A 662, A18 (2022).
@@
- Militzer et al. [2022] B. Militzer, W. B. Hubbard, S. Wahl, J. I. Lunine, E. Galanti, Y. Kaspi, Y. Miguel, T. Guillot, K. M. Moore, M. Parisi, J. E. P. Connerney, R. Helled, H. Cao, C. Mankovich, D. J. Stevenson, R. S. Park, M. Wong, S. K. Atreya, J. Anderson, and S. J. Bolton, Juno Spacecraft Measurements of Jupiter’s Gravity Imply a Dilute Core, The Planetary Science Journal 3, 185 (2022).
@@
- Chabrier et al. [2019] G. Chabrier, S. Mazevet, and F. Soubiran, A new equation of state for dense hydrogen–helium mixtures, The Astrophysical Journal 872, 51 (2019).
@@
- Militzer and Hubbard [2013] B. Militzer and W. B. Hubbard, Ab initio equation of state for hydrogen–helium mixtures with recalibration of the giant-planet mass–radius relation, The Astrophysical Journal 774, 148 (2013).
@@
- Nettelmann et al. [2012] N. Nettelmann, A. Becker, B. Holst, and R. Redmer, Jupiter models with improved ab initio hydrogen equation of state (h-reos.2), The Astrophysical Journal 750, 52 (2012).
@@
- Becker et al. [2014] A. Becker, W. Lorenzen, J. J. Fortney, N. Nettelmann, M. Schöttler, and R. Redmer, Ab initio equations of state for hydrogen (h-reos.3) and helium (he-reos.3) and their implications for the interior of brown dwarfs, The Astrophysical Journal Supplement Series 215, 21 (2014).
@@
- Perdew et al. [1996] J. P. Perdew, K. Burke, and M. Ernzerhof, Generalized gradient approximation made simple, Phys. Rev. Lett. 77, 3865 (1996).
@@
- Kirkwood [1935] J. G. Kirkwood, Statistical Mechanics of Fluid Mixtures, The Journal of Chemical Physics 3, 300 (1935).
@@
- de Wijs et al. [1998] G. A. de Wijs, G. Kresse, and M. J. Gillan, First-order phase transitions by first-principles free-energy calculations: The melting of al, Phys. Rev. B 57, 8223 (1998).
@@
- Frenkel and Smit [2023] D. Frenkel and B. Smit, Understanding Molecular Simulation: From Algorithms to Applications, 3rd ed. (Academic Press, 2023).
@@
- Morales et al. [2010a] M. A. Morales, C. Pierleoni, and D. M. Ceperley, Equation of state of metallic hydrogen from coupled electron-ion monte carlo simulations, Phys. Rev. E 81, 021202 (2010a).
@@
- Bergermann et al. [2024] A. Bergermann, M. French, and R. Redmer, Ab initio calculation of the miscibility diagram for mixtures of hydrogen and water, Phys. Rev. B 109, 174107 (2024).
@@
- Saumon et al. [1995] D. Saumon, G. Chabrier, and H. M. van Horn, An Equation of State for Low-Mass Stars and Giant Planets, Astrophysical Journal Supplement 99, 713 (1995).
@@
- Miguel, Y. et al. [2018] Miguel, Y., Guillot, T., and Fayon, L., Jupiter internal structure: the effect of different equations of state (corrigendum), A&A 618, C2 (2018).
@@
- Albergo and Vanden-Eijnden [2023] M. S. Albergo and E. Vanden-Eijnden, Building normalizing flows with stochastic interpolants, in The Eleventh International Conference on Learning Representations (2023).
@@
- Lipman et al. [2023] Y. Lipman, R. T. Q. Chen, H. Ben-Hamu, M. Nickel, and M. Le, Flow matching for generative modeling, in The Eleventh International Conference on Learning Representations (2023).
@@
- Liu et al. [2023] X. Liu, C. Gong, and qiang liu, Flow straight and fast: Learning to generate and transfer data with rectified flow, in The Eleventh International Conference on Learning Representations (2023).
- Zhao and Wang [2023] L. Zhao and L. Wang, Bounding free energy difference with flow matching, Chinese Physics Letters 40, 120201 (2023).
- [27] See Supplemental Material for more details on the theoretical formulations of targeted free energy perturbation and flow matching, as well as their implementation and some illustrative examples in practice. There are also further discussions about the effects of thermodynamic inconsistency on entropy calculation, our new EOS construction approach, and the comparison between Jupiter’s adiabats derived from various hydrogen EOSs.
- Jarzynski [2002] C. Jarzynski, Targeted free energy perturbation, Phys. Rev. E 65, 046122 (2002).
@@
- Hahn and Then [2009] A. M. Hahn and H. Then, Using bijective maps to improve free-energy estimates, Phys. Rev. E 79, 011113 (2009).
@@
- Papamakarios et al. [2021] G. Papamakarios, E. Nalisnick, D. J. Rezende, S. Mohamed, and B. Lakshminarayanan, Normalizing flows for probabilistic modeling and inference, Journal of Machine Learning Research 22, 1 (2021).
- Wirnsberger et al. [2020] P. Wirnsberger, A. J. Ballard, G. Papamakarios, S. Abercrombie, S. Racanière, A. Pritzel, D. Jimenez Rezende, and C. Blundell, Targeted free energy estimation via learned mappings, The Journal of Chemical Physics 153, 144112 (2020).
- Wirnsberger et al. [2022] P. Wirnsberger, G. Papamakarios, B. Ibarz, S. Racanière, A. J. Ballard, A. Pritzel, and C. Blundell, Normalizing flows for atomic solids, Machine Learning: Science and Technology 3, 025009 (2022).
- Ahmad and Cai [2022] R. Ahmad and W. Cai, Free energy calculation of crystalline solids using normalizing flows, Modelling and Simulation in Materials Science and Engineering 30, 065007 (2022).
- Molina-Taborda et al. [2024] A. Molina-Taborda, P. Cossio, O. Lopez-Acevedo, and M. Gabrié, Active Learning of Boltzmann Samplers and Potential Energies with Quantum Mechanical Accuracy, Journal of Chemical Theory and Computation 20, 8833 (2024).
- Li and Wang [2018] S.-H. Li and L. Wang, Neural network renormalization group, Phys. Rev. Lett. 121, 260601 (2018).
- Wu et al. [2019] D. Wu, L. Wang, and P. Zhang, Solving statistical mechanics using variational autoregressive networks, Phys. Rev. Lett. 122, 080602 (2019).
- Xie et al. [2022] H. Xie, L. Zhang, and L. Wang, Ab-initio study of interacting fermions at finite temperature with neural canonical transformation, Journal of Machine Learning 1, 38 (2022).
- Xie et al. [2023a] H. Xie, L. Zhang, and L. Wang, of two-dimensional electron gas: A neural canonical transformation study, SciPost Phys. 14, 154 (2023a).
- Xie et al. [2023b] H. Xie, Z.-H. Li, H. Wang, L. Zhang, and L. Wang, Deep variational free energy approach to dense hydrogen, Phys. Rev. Lett. 131, 126501 (2023b).
@@
- Lorenzen et al. [2010] W. Lorenzen, B. Holst, and R. Redmer, First-order liquid-liquid phase transition in dense hydrogen, Phys. Rev. B 82, 195107 (2010).
@@
- Geng et al. [2019] H. Y. Geng, Q. Wu, M. Marqués, and G. J. Ackland, Thermodynamic anomalies and three distinct liquid-liquid transitions in warm dense liquid hydrogen, Phys. Rev. B 100, 134109 (2019).
@@
- Caillabet et al. [2011] L. Caillabet, S. Mazevet, and P. Loubeyre, Multiphase equation of state of hydrogen from ab initio calculations in the range 0.2 to 5 g/cc up to 10 ev, Phys. Rev. B 83, 094101 (2011).
@@
- Holst et al. [2008] B. Holst, R. Redmer, and M. P. Desjarlais, Thermophysical properties of warm dense hydrogen using quantum molecular dynamics simulations, Phys. Rev. B 77, 184201 (2008).
@@
- Juranek et al. [2002] H. Juranek, R. Redmer, and Y. Rosenfeld, Fluid variational theory for pressure dissociation in dense hydrogen: Multicomponent reference system and nonadditivity effects, The Journal of Chemical Physics 117, 1768 (2002).
@@
- Holst et al. [2007] B. Holst, N. Nettelmann, and R. Redmer, Equation of state for dense hydrogen and plasma phase transition, Contributions to Plasma Physics 47, 368 (2007).
@@
- Cox et al. [1989] J. D. Cox, D. D. Wagman, and V. A. Medvedev, Codata key values for thermodynamics, in Series on Thermodynamic Properties (1989).
@@
- Lemmon et al. [2023] E. W. Lemmon, I. H. Bell, M. L. Huber, and M. O. McLinden, Thermophysical properties of fluid systems, in NIST Chemistry WebBook, NIST Standard Reference Database Number 69, edited by P. Linstrom and W. Mallard (National Institute of Standards and Technology, Gaithersburg MD, 20899, 2023).
- NIS [2022] NIST Computational Chemistry Comparison and Benchmark Database, NIST Standard Reference Database Number 101 (2022).
@@
- Schöttler and Redmer [2018] M. Schöttler and R. Redmer, Ab initio calculation of the miscibility diagram for hydrogen-helium mixtures, Phys. Rev. Lett. 120, 115703 (2018).
@@
- Vorberger et al. [2007] J. Vorberger, I. Tamblyn, B. Militzer, and S. A. Bonev, Hydrogen-helium mixtures in the interiors of giant planets, Phys. Rev. B 75, 024206 (2007).
@@
- Chang et al. [2024] X. Chang, B. Chen, Q. Zeng, H. Wang, K. Chen, Q. Tong, X. Yu, D. Kang, S. Zhang, F. Guo, Y. Hou, Z. Zhao, Y. Yao, Y. Ma, and J. Dai, Theoretical evidence of h-he demixing under jupiter and saturn conditions, Nature Communications 15, 8543 (2024).
@@
- Pierleoni et al. [2016] C. Pierleoni, M. A. Morales, G. Rillo, M. Holzmann, and D. M. Ceperley, Liquid–liquid phase transition in hydrogen by coupled electron–ion monte carlo simulations, Proceedings of the National Academy of Sciences 113, 4953 (2016).
@@
- Mazzola et al. [2018] G. Mazzola, R. Helled, and S. Sorella, Phase diagram of hydrogen and a hydrogen-helium mixture at planetary conditions by quantum monte carlo simulations, Phys. Rev. Lett. 120, 025701 (2018).
@@
- Nakano et al. [2020] K. Nakano, C. Attaccalite, M. Barborini, L. Capriotti, M. Casula, E. Coccia, M. Dagrada, C. Genovese, Y. Luo, G. Mazzola, A. Zen, and S. Sorella, TurboRVB: A many-body toolkit for ab initio electronic simulations by quantum Monte Carlo, The Journal of Chemical Physics 152, 204121 (2020).
@@
- Morales et al. [2010b] M. A. Morales, C. Pierleoni, E. Schwegler, and D. M. Ceperley, Evidence for a first-order liquid-liquid transition in high-pressure hydrogen from ab initio simulations, Proceedings of the National Academy of Sciences 107, 12799 (2010b).
@@
- Niu et al. [2023] H. Niu, Y. Yang, S. Jensen, M. Holzmann, C. Pierleoni, and D. M. Ceperley, Stable solid molecular hydrogen above 900 k from a machine-learned potential trained with diffusion quantum monte carlo, Phys. Rev. Lett. 130, 076102 (2023).
@@
- Wang et al. [2023] X. Wang, Z. Wang, P. Gao, C. Zhang, J. Lv, H. Wang, H. Liu, Y. Wang, and Y. Ma, Data-driven prediction of complex crystal structures of dense lithium, Nature Communications 14, 2924 (2023).
@@
- Zwanzig [1954] R. W. Zwanzig, High‐Temperature Equation of State by a Perturbation Method. I. Nonpolar Gases, The Journal of Chemical Physics 22, 1420 (1954).
@@
- Bennett [1976] C. H. Bennett, Efficient estimation of free energy differences from monte carlo data, Journal of Computational Physics 22, 245 (1976).
@@
- Chen et al. [2018] R. T. Q. Chen, Y. Rubanova, J. Bettencourt, and D. Duvenaud, Neural ordinary differential equations, in Proceedings of the 32nd International Conference on Neural Information Processing Systems, NIPS’18 (Curran Associates Inc., Red Hook, NY, USA, 2018) p. 6572–6583.
@@
- Zhang et al. [2018] L. Zhang, W. E, and L. Wang, Monge-ampère flow for generative modeling (2018), arXiv:1809.10188 [cs.LG] .
@@
- Vaswani et al. [2017] A. Vaswani, N. Shazeer, N. Parmar, J. Uszkoreit, L. Jones, A. N. Gomez, L. u. Kaiser, and I. Polosukhin, Attention is all you need, in Advances in Neural Information Processing Systems, Vol. 30, edited by I. Guyon, U. V. Luxburg, S. Bengio, H. Wallach, R. Fergus, S. Vishwanathan, and R. Garnett (Curran Associates, Inc., 2017).
@@
- Köhler et al. [2019] J. Köhler, L. Klein, and F. Noé, Equivariant flows: sampling configurations for multi-body systems with symmetric energies (2019), arXiv:1910.00753 [stat.ML] .

### v2

@@ Acknowledgements.
We acknowledge useful discussions with Lei Wang, Ravit Helled, Didier Saumon, Gilles Chabrier, and Armin Bergermann. We also acknowledge Cesare Cozza for providing initial data and support with the DFT calculations. H.X. and G.M. acknowledge financial support from the Swiss National Science Foundation (Grant No. PCEFP2_203455). S.H. also acknowledges financial support from the Swiss National Science Foundation (Grant No. 200020_215634).References
@@
- Helled et al. (2020) Ravit Helled, Guglielmo Mazzola, and Ronald Redmer, “Understanding dense hydrogen at planetary conditions,” Nature Reviews Physics 2, 562–574 (2020).
@@
- Johnson et al. (1992) T. V. Johnson, C. M. Yeates, and R. Young, “Space science reviews volume on Galileo Mission overview,” Space Science Reviews 60, 3–21 (1992).
@@
- Bolton et al. (2017) S. J. Bolton, J. Lunine, D. Stevenson, J. E. P. Connerney, S. Levin, T. C. Owen, F. Bagenal, D. Gautier, A. P. Ingersoll, G. S. Orton, T. Guillot, W. Hubbard, J. Bloxham, A. Coradini, S. K. Stephens, P. Mokashi, R. Thorne, and R. Thorpe, “The Juno Mission,” Space Science Reviews 213, 5–37 (2017).
@@
- Miguel, Y. et al. (2016) Miguel, Y., Guillot, T., and Fayon, L., “Jupiter internal structure: the effect of different equations of state,” A&A 596, A114 (2016).
@@
- Debras and Chabrier (2019) Florian Debras and Gilles Chabrier, “New Models of Jupiter in the Context of Juno and Galileo,” The Astrophysical Journal 872, 100 (2019).
@@
- Mazevet, S. et al. (2022) Mazevet, S., Licari, A., and Soubiran, F., “Benchmarking the ab initio hydrogen equation of state for the interior structure of jupiter,” A&A 664, A112 (2022).
@@
- Howard, S. et al. (2023) Howard, S., Guillot, T., Bazot, M., Miguel, Y., Stevenson, D. J., Galanti, E., Kaspi, Y., Hubbard, W. B., Militzer, B., Helled, R., Nettelmann, N., Idini, B., and Bolton, S., “Jupiter’s interior from Juno: Equation-of-state uncertainties and dilute core extent,” A&A 672, A33 (2023).
@@
- Nettelmann et al. (2021) N. Nettelmann, N. Movshovitz, D. Ni, J. J. Fortney, E. Galanti, Y. Kaspi, R. Helled, C. R. Mankovich, and S. Bolton, “Theory of Figures to the Seventh Order and the Interiors of Jupiter and Saturn,” The Planetary Science Journal 2, 241 (2021).
@@
- Miguel, Y. et al. (2022) Miguel, Y., Bazot, M., Guillot, T., Howard, S., Galanti, E., Kaspi, Y., Hubbard, W. B., Militzer, B., Helled, R., Atreya, S. K., Connerney, J. E. P., Durante, D., Kulowski, L., Lunine, J. I., Stevenson, D., and Bolton, S., “Jupiter’s inhomogeneous envelope,” A&A 662, A18 (2022).
@@
- Militzer et al. (2022) Burkhard Militzer, William B. Hubbard, Sean Wahl, Jonathan I. Lunine, Eli Galanti, Yohai Kaspi, Yamila Miguel, Tristan Guillot, Kimberly M. Moore, Marzia Parisi, John E. P. Connerney, Ravid Helled, Hao Cao, Christopher Mankovich, David J. Stevenson, Ryan S. Park, Mike Wong, Sushil K. Atreya, John Anderson, and Scott J. Bolton, “Juno Spacecraft Measurements of Jupiter’s Gravity Imply a Dilute Core,” The Planetary Science Journal 3, 185 (2022).
@@
- Chabrier et al. (2019) G. Chabrier, S. Mazevet, and F. Soubiran, “A new equation of state for dense hydrogen–helium mixtures,” The Astrophysical Journal 872, 51 (2019).
@@
- Militzer and Hubbard (2013) B. Militzer and W. B. Hubbard, “Ab initio equation of state for hydrogen–helium mixtures with recalibration of the giant-planet mass–radius relation,” The Astrophysical Journal 774, 148 (2013).
@@
- Nettelmann et al. (2012) N. Nettelmann, A. Becker, B. Holst, and R. Redmer, “Jupiter models with improved ab initio hydrogen equation of state (h-reos.2),” The Astrophysical Journal 750, 52 (2012).
@@
- Becker et al. (2014) Andreas Becker, Winfried Lorenzen, Jonathan J. Fortney, Nadine Nettelmann, Manuel Schöttler, and Ronald Redmer, “Ab initio equations of state for hydrogen (h-reos.3) and helium (he-reos.3) and their implications for the interior of brown dwarfs,” The Astrophysical Journal Supplement Series 215, 21 (2014).
@@
- Perdew et al. (1996) John P. Perdew, Kieron Burke, and Matthias Ernzerhof, “Generalized gradient approximation made simple,” Phys. Rev. Lett. 77, 3865–3868 (1996).
@@
- Kirkwood (1935) John G. Kirkwood, “Statistical Mechanics of Fluid Mixtures,” The Journal of Chemical Physics 3, 300–313 (1935).
@@
- de Wijs et al. (1998) Gilles A. de Wijs, Georg Kresse, and Michael J. Gillan, “First-order phase transitions by first-principles free-energy calculations: The melting of al,” Phys. Rev. B 57, 8223–8234 (1998).
@@
- Frenkel and Smit (2023) D. Frenkel and B. Smit, Understanding Molecular Simulation: From Algorithms to Applications, 3rd ed. (Academic Press, 2023).
@@
- Morales et al. (2010a) Miguel A. Morales, Carlo Pierleoni, and D. M. Ceperley, “Equation of state of metallic hydrogen from coupled electron-ion monte carlo simulations,” Phys. Rev. E 81, 021202 (2010a).
@@
- Bergermann et al. (2024) Armin Bergermann, Martin French, and Ronald Redmer, “Ab initio calculation of the miscibility diagram for mixtures of hydrogen and water,” Phys. Rev. B 109, 174107 (2024).
@@
- Saumon et al. (1995) D. Saumon, G. Chabrier, and H. M. van Horn, “An Equation of State for Low-Mass Stars and Giant Planets,” Astrophysical Journal Supplement 99, 713 (1995).
@@
- Miguel, Y. et al. (2018) Miguel, Y., Guillot, T., and Fayon, L., “Jupiter internal structure: the effect of different equations of state (corrigendum),” A&A 618, C2 (2018).
@@
- Albergo and Vanden-Eijnden (2023) Michael Samuel Albergo and Eric Vanden-Eijnden, “Building normalizing flows with stochastic interpolants,” in The Eleventh International Conference on Learning Representations (2023).
@@
- Lipman et al. (2023) Yaron Lipman, Ricky T. Q. Chen, Heli Ben-Hamu, Maximilian Nickel, and Matthew Le, “Flow matching for generative modeling,” in The Eleventh International Conference on Learning Representations (2023).
@@
- Liu et al. (2023) Xingchao Liu, Chengyue Gong, and qiang liu, “Flow straight and fast: Learning to generate and transfer data with rectified flow,” in The Eleventh International Conference on Learning Representations (2023).
- Jarzynski (2002) C. Jarzynski, “Targeted free energy perturbation,” Phys. Rev. E 65, 046122 (2002).
@@
- Hahn and Then (2009) A. M. Hahn and H. Then, “Using bijective maps to improve free-energy estimates,” Phys. Rev. E 79, 011113 (2009).
- (28) See Supplemental Material for more details on the theoretical formulations of targeted free energy perturbation and flow matching, as well as their implementation and some illustrative examples in practice. There are also further discussions about the effects of thermodynamic inconsistency on entropy calculation, our new EOS construction approach, and the comparison between Jupiter’s adiabats derived from various hydrogen EOSs.
@@
- Papamakarios et al. (2021) George Papamakarios, Eric Nalisnick, Danilo Jimenez Rezende, Shakir Mohamed, and Balaji Lakshminarayanan, “Normalizing flows for probabilistic modeling and inference,” Journal of Machine Learning Research 22, 1–64 (2021).
- Zhao and Wang (2023) Lu Zhao and Lei Wang, “Bounding free energy difference with flow matching,” Chinese Physics Letters 40, 120201 (2023).
@@
- Lorenzen et al. (2010) Winfried Lorenzen, Bastian Holst, and Ronald Redmer, “First-order liquid-liquid phase transition in dense hydrogen,” Phys. Rev. B 82, 195107 (2010).
@@
- Geng et al. (2019) Hua Y. Geng, Q. Wu, Miriam Marqués, and Graeme J. Ackland, “Thermodynamic anomalies and three distinct liquid-liquid transitions in warm dense liquid hydrogen,” Phys. Rev. B 100, 134109 (2019).
@@
- Caillabet et al. (2011) L. Caillabet, S. Mazevet, and P. Loubeyre, “Multiphase equation of state of hydrogen from ab initio calculations in the range 0.2 to 5 g/cc up to 10 ev,” Phys. Rev. B 83, 094101 (2011).
@@
- Holst et al. (2008) Bastian Holst, Ronald Redmer, and Michael P. Desjarlais, “Thermophysical properties of warm dense hydrogen using quantum molecular dynamics simulations,” Phys. Rev. B 77, 184201 (2008).
@@
- Juranek et al. (2002) Hauke Juranek, Ronald Redmer, and Yaakov Rosenfeld, “Fluid variational theory for pressure dissociation in dense hydrogen: Multicomponent reference system and nonadditivity effects,” The Journal of Chemical Physics 117, 1768–1774 (2002).
@@
- Holst et al. (2007) B. Holst, N. Nettelmann, and R. Redmer, “Equation of state for dense hydrogen and plasma phase transition,” Contributions to Plasma Physics 47, 368–374 (2007).
@@
- Cox et al. (1989) John D Cox, Donald D Wagman, and Vadim Andreevich Medvedev, “Codata key values for thermodynamics,” in Series on Thermodynamic Properties (1989).
@@
- Lemmon et al. (2023) Eric W. Lemmon, Ian H. Bell, Marcia L. Huber, and Mark O. McLinden, “Thermophysical properties of fluid systems,” in NIST Chemistry WebBook, NIST Standard Reference Database Number 69, edited by P.J. Linstrom and W.G. Mallard (National Institute of Standards and Technology, Gaithersburg MD, 20899, 2023).
- (39) Didier Saumon and Gilles Chabrier, private communication.
@@
- Schöttler and Redmer (2018) Manuel Schöttler and Ronald Redmer, “Ab initio calculation of the miscibility diagram for hydrogen-helium mixtures,” Phys. Rev. Lett. 120, 115703 (2018).
@@
- Vorberger et al. (2007) J. Vorberger, I. Tamblyn, B. Militzer, and S. A. Bonev, “Hydrogen-helium mixtures in the interiors of giant planets,” Phys. Rev. B 75, 024206 (2007).
@@
- Chang et al. (2024) Xiaoju Chang, Bo Chen, Qiyu Zeng, Han Wang, Kaiguo Chen, Qunchao Tong, Xiaoxiang Yu, Dongdong Kang, Shen Zhang, Fangyu Guo, Yong Hou, Zengxiu Zhao, Yansun Yao, Yanming Ma, and Jiayu Dai, “Theoretical evidence of h-he demixing under jupiter and saturn conditions,” Nature Communications 15, 8543 (2024).
@@
- Pierleoni et al. (2016) Carlo Pierleoni, Miguel A. Morales, Giovanni Rillo, Markus Holzmann, and David M. Ceperley, “Liquid–liquid phase transition in hydrogen by coupled electron–ion monte carlo simulations,” Proceedings of the National Academy of Sciences 113, 4953–4957 (2016).
@@
- Mazzola et al. (2018) Guglielmo Mazzola, Ravit Helled, and Sandro Sorella, “Phase diagram of hydrogen and a hydrogen-helium mixture at planetary conditions by quantum monte carlo simulations,” Phys. Rev. Lett. 120, 025701 (2018).
@@
- Nakano et al. (2020) Kousuke Nakano, Claudio Attaccalite, Matteo Barborini, Luca Capriotti, Michele Casula, Emanuele Coccia, Mario Dagrada, Claudio Genovese, Ye Luo, Guglielmo Mazzola, Andrea Zen, and Sandro Sorella, “TurboRVB: A many-body toolkit for ab initio electronic simulations by quantum Monte Carlo,” The Journal of Chemical Physics 152, 204121 (2020).
- Xie et al. (2023a) Hao Xie, Zi-Hang Li, Han Wang, Linfeng Zhang, and Lei Wang, “Deep variational free energy approach to dense hydrogen,” Phys. Rev. Lett. 131, 126501 (2023a).
@@
- Morales et al. (2010b) Miguel A. Morales, Carlo Pierleoni, Eric Schwegler, and D. M. Ceperley, “Evidence for a first-order liquid-liquid transition in high-pressure hydrogen from ab initio simulations,” Proceedings of the National Academy of Sciences 107, 12799–12803 (2010b).
@@
- Niu et al. (2023) Hongwei Niu, Yubo Yang, Scott Jensen, Markus Holzmann, Carlo Pierleoni, and David M. Ceperley, “Stable solid molecular hydrogen above 900 k from a machine-learned potential trained with diffusion quantum monte carlo,” Phys. Rev. Lett. 130, 076102 (2023).
@@
- Wang et al. (2023) Xiaoyang Wang, Zhenyu Wang, Pengyue Gao, Chengqian Zhang, Jian Lv, Han Wang, Haifeng Liu, Yanchao Wang, and Yanming Ma, “Data-driven prediction of complex crystal structures of dense lithium,” Nature Communications 14, 2924 (2023).
@@
- Zwanzig (1954) Robert W. Zwanzig, “High‐Temperature Equation of State by a Perturbation Method. I. Nonpolar Gases,” The Journal of Chemical Physics 22, 1420–1426 (1954).
@@
- Bennett (1976) Charles H Bennett, “Efficient estimation of free energy differences from monte carlo data,” Journal of Computational Physics 22, 245–268 (1976).
- Wirnsberger et al. (2020) Peter Wirnsberger, Andrew J. Ballard, George Papamakarios, Stuart Abercrombie, Sébastien Racanière, Alexander Pritzel, Danilo Jimenez Rezende, and Charles Blundell, “Targeted free energy estimation via learned mappings,” The Journal of Chemical Physics 153, 144112 (2020).
- Wirnsberger et al. (2022) Peter Wirnsberger, George Papamakarios, Borja Ibarz, Sébastien Racanière, Andrew J Ballard, Alexander Pritzel, and Charles Blundell, “Normalizing flows for atomic solids,” Machine Learning: Science and Technology 3, 025009 (2022).
- Ahmad and Cai (2022) Rasool Ahmad and Wei Cai, “Free energy calculation of crystalline solids using normalizing flows,” Modelling and Simulation in Materials Science and Engineering 30, 065007 (2022).
- Molina-Taborda et al. (2024) Ana Molina-Taborda, Pilar Cossio, Olga Lopez-Acevedo, and Marylou Gabrié, “Active Learning of Boltzmann Samplers and Potential Energies with Quantum Mechanical Accuracy,” Journal of Chemical Theory and Computation 20, 8833–8843 (2024).
- Li and Wang (2018) Shuo-Hui Li and Lei Wang, “Neural network renormalization group,” Phys. Rev. Lett. 121, 260601 (2018).
- Wu et al. (2019) Dian Wu, Lei Wang, and Pan Zhang, “Solving statistical mechanics using variational autoregressive networks,” Phys. Rev. Lett. 122, 080602 (2019).
- Xie et al. (2022) Hao Xie, Linfeng Zhang, and Lei Wang, “Ab-initio study of interacting fermions at finite temperature with neural canonical transformation,” Journal of Machine Learning 1, 38–59 (2022).
- Xie et al. (2023b) Hao Xie, Linfeng Zhang, and Lei Wang, “ of two-dimensional electron gas: A neural canonical transformation study,” SciPost Phys. 14, 154 (2023b).
@@
- Chen et al. (2018) Ricky T. Q. Chen, Yulia Rubanova, Jesse Bettencourt, and David Duvenaud, “Neural ordinary differential equations,” in Proceedings of the 32nd International Conference on Neural Information Processing Systems, NIPS’18 (Curran Associates Inc., Red Hook, NY, USA, 2018) p. 6572–6583.
@@
- Zhang et al. (2018) Linfeng Zhang, Weinan E, and Lei Wang, “Monge-ampère flow for generative modeling,” (2018), arXiv:1809.10188 [cs.LG] .
@@
- Vaswani et al. (2017) Ashish Vaswani, Noam Shazeer, Niki Parmar, Jakob Uszkoreit, Llion Jones, Aidan N Gomez, Ł ukasz Kaiser, and Illia Polosukhin, “Attention is all you need,” in Advances in Neural Information Processing Systems, Vol. 30, edited by I. Guyon, U. Von Luxburg, S. Bengio, H. Wallach, R. Fergus, S. Vishwanathan, and R. Garnett (Curran Associates, Inc., 2017).
@@
- Köhler et al. (2019) Jonas Köhler, Leon Klein, and Frank Noé, “Equivariant flows: sampling configurations for multi-body systems with symmetric energies,” (2019), arXiv:1910.00753 [stat.ML] .